# A model-independent measurement of the expansion and growth rates from BOSS using the FreePower method

### v1

@@ 2 The FreePower model-independent approach
Previous tests indicate that neither the exact spacing nor the number of the $k$-bins has a strong influence on the results on the other parameters [38]. The linear matter power spectrum is constructed with these varied parameters and a cubic spline interpolation between them. Because of that, the linear matter power spectrum is quite smooth. Therefore, the coupling between the large-scale IR modes and the mildly non-linear modes is rather weak [10] and the effect of IR resummation is in general small. In this approach we do not disentangle the growth factor $G(z)$ from the linear power spectrum, so we in fact measure $P_{L}(k,z)=G(z)^{2}P_{L}(k)$ at each redshift bin.
@@ 2.2 Data sets from BOSS
linear growth rate $f$ ${\cal U}[0,1]$ ${\cal U}[0,1]$
@@
dim.less angular diameter distance $L_{A}$ $\mathcal{N}(0.25,0.008)$ $\mathcal{N}(0.323,0.010)$
linear bias $b_{1}$ $\mathcal{N}(1.67,0.36)$ ${\cal U}[0,4]$
${\cal G}_{2}$ bias $b_{\mathcal{G}_{2}}$ $\mathcal{N}(0.11,0.68)$ $\mathcal{N}(0,1)$
quadratic bias $b_{2}$ $\mathcal{N}(-2.3,1.4)$ $\mathcal{N}(-3,2)$
@@
counterterm (monopole) $c_{0}/[{\rm Mpc}/h]^{2}$ $\mathcal{N}(-18.5,95.7)$ $\mathcal{N}(0,30)$
counterterm (quadrupole) $c_{2}/[{\rm Mpc}/h]^{2}$ $\mathcal{N}(14.7,72)$ $\mathcal{N}(30,30)$
@@
$P_{7}/$$[{\rm Mpc}/h]^{3}$ $\mathcal{N}(375,50)$ $\mathcal{N}(375,500)$
$P_{8}/$$[{\rm Mpc}/h]^{3}$ $\mathcal{N}(100,20)$ $\mathcal{N}(100,200)$
@@
For the bias parameters, we left the priors very broad for high-$z$. For low-$z$, in order to see if convergence of the chain could be further improved, we used bias priors which were a bit tighter, while at the same time still broader than the likelihoods. We used as reference the final NGC low-$z$ posterior results from ISZ20. We took their errorbars, propagated the uncertainty in the power spectrum amplitude $A$ since they multiply all their bias parameters by $(A/A_{\rm Planck})^{1/2}$, symmetrized the results and multiplied them by a factor of two. We used the final results for the mean and standard deviation to construct our low-$z$ uncorrelated Gaussian priors. We also did the same for the counterterm parameters. We thus allow a twice larger uncertainty of ISZ20 in each bias and counterterm parameter as our prior. We also chose to adopt $10$ times tighter priors for the very small-scale bins $P_{7}$ and $P_{8}$ since these are beyond our 1-loop range and thus carry little information on $f$ and $E$. The Gaussian prior on $L_{A}(z)$ is centered around its Planck 2018 $\Lambda$CDM value (i.e. assuming $\Omega_{m0}=0.31$ so that $L_{A}=0.25$ for the low-$z$ and 0.323 for the high-$z$ sample) with a $3\%$ standard deviation. For $z\leq 1$ current constraints from Pantheon+ [46] are already at this level.
@@ 3 MCMC results
We ran emcee using 256 walkers initialized in a small volume in parameter space. We ran, for each walker, 270k (180k) steps, for a total of 69 (46) million chain points for the low-$z$ (high-$z$) samples. Due to the complexity of the 1-loop calculations and the amount of MCMC steps needed, the total computational time used for both redshift bin runs was large, around 70000 CPU-hours. We analyzed the results using both the Python package getdist [47] and our own codes. The extra 50% points used in low-$z$ was to confirm that the results we found, which as we will discuss is in tension with $\Lambda$CDM CMB results, were robust. Indeed, the results adding these extra points changed little, which gives us confidence also in the convergence of our high-$z$ results.
@@
Tables 2 summarizes the marginalized constraints in each cosmological and bias parameter, whereas Table 3 show the results for all 8 $k$-bins of $P(k)$. Both tables also show the best-fit Planck 2018 values assuming flat $\Lambda$CDM.
The correlations $\rho$ among the 17 parameters are in most cases small. The average off-diagonal absolute value of $\rho$ is around 0.19 for both redshift bins. We list here some particular values for $z_{\rm eff}=0.38$ (0.61): $\rho(f,E)=0.59$ ($0.74$), $\rho(E,L_{A})=-0.28$ ($-0.27$), $\rho(E,b_{1})=-0.53$ ($-0.78$), $\rho(E,b_{\mathcal{G}_{2}})=-0.66$ ($-0.50$), $\rho(b_{\mathcal{G}_{2}},c_{2})=-0.79$ ($-0.19$). The full correlation matrices for both bins are shown in Appendix A.
One could worry that the informative prior we imposed on some of the parameters might have biased the low-$z$ results. Let’s dig deeper into this issue. First, the anti-correlation between $E,\,b_{1}$ (see Figures 6 and 8 in the appendices) means that a broader prior for $b_{1}$ could only allow larger values of $b_{1}$ and thereby increase the tension of $E$ with $\Lambda$CDM. The same holds true for $b_{{\cal G}_{2}}$, although the prior is less informative here. For $b_{2}$ the prior is the most informative and the correlation is positive ($\rho=0.33$), so a broader prior could allow for larger $b_{2}$, and in this case could indeed induce a higher $E$, alleviating the tension. However, the correlation is mild and our result for $b_{2}$ is in agreement with previous analyses of BOSS data [14, 17], so we do not expect that a broader prior would significantly affect the final constraints on $E$.
@@
$f$ $0.60^{+0.11}_{-0.09}$ $0.69^{+0.14}_{-0.13}$ 0.716 0.792
$E$ $1.052^{+0.060}_{-0.041}$ $1.36\pm{0.10}$ 1.23 1.41
$L_{A}$ $0.248^{+0.008}_{-0.007}$ $0.323^{+0.009}_{-0.010}$ 0.250 0.323
$b_{1}$ $2.08^{+0.12}_{-0.13}$ $2.03\pm 0.17$
$b_{2}$ $-2.4^{+1.1}_{-1.0}$ $-2.0^{+1.1}_{-1.0}$
$b_{\mathcal{G}_{2}}$ $0.33^{+0.28}_{-0.25}$ $0.07^{+0.31}_{-0.36}$
$c_{0}/[{\rm Mpc}/h]^{2}$ $-10^{+26}_{-36}$ $12\pm 18$
$c_{2}/[{\rm Mpc}/h]^{2}$ $27^{+23}_{-27}$ $26^{+21}_{-19}$
$s$ $0.26^{+0.48}_{-0.45}$ $0.14^{+0.43}_{-0.42}$
Table 2: Main results for the two redshift bins. We show best-fits with 68.3% CI. The last column shows the expected values for $\Lambda$CDM with Planck 2018 values [48].
@@
Table 3: Best-fit and 68.3% CI for each $k$-bin of the linear power spectrum (at $z=0$), in units of $({\rm Mpc}/h)^{3}$. The $k$ values represent the center of each $k$-bin. The model-independent constraints on $P_{i}(z)$ were converted to $z=0$ values for comparison using the $\Lambda$CDM growth factor $G(z)=0.818\hskip 5.69054pt(0.728)$ for $z=0.38\hskip 5.69054pt(0.61)$. The last row reports the corresponding Planck 2018 $\Lambda$CDM values as reference values.
We tested the convergence of our MCMC using primarly the auto-correlation time ($\tau$) estimates [37]. A large ratio between the number of steps and $\tau$ indicate that the chain ran for long enough to have a large effective number of independent steps. Ideally $\tau$ should converge to an asymptotic value for high enough number of steps. We found that around 200000 steps with each of 256 walkers were needed for reaching convergence. More details are provided in Appendix A. We also performed the Gelman-Rubin test [49]. This simple test is not the most reliable, specially using an ensemble sampler like emcee since in this approach the walkers are not independent. Nevertheless it gives another handle on convergence which is independent of the auto-correlation time analysis. For the cosmological parameters $f$, $E$ and $L_{A}$, we get $(\hat{R})^{(1/2)}=\{1.012,1.019,1.011\}$ for the low-$z$ run, and $(\hat{R})^{(1/2)}=\{1.071,1.091,1.027\}$ for the high-$z$ run. It is commonly assumed as a threshold for convergence the value $(\hat{R}_{\rm max})^{(1/2)}=1.1$. This agrees with the auto-correlation time results, which indicate good (reasonable) convergence for the low-$z$ (high-$z$) chains.
@@ 4 Comparison with previous analyses of BOSS data
A comparison of the inferred cosmological parameters with these works is given in Table 4. In this Table we quote mean values and 68% percentiles using equal-tailed statistics, instead of highest-density values, since those are the ones reported in ISZ20 and PI22. For the low-$z$ survey we find $f=0.62\pm 0.10$ and $E=1.07\pm 0.05$, in good agreement with $f=0.74\pm 0.08,E=1.14\pm 0.05$ from ISZ20. FreePower measures directly $f(z)$, while conventional analysis measure instead $f(z)\sigma_{8}(z)$. We therefore convert the latter using the values of $\sigma_{8}$ and $\Omega_{m0}$ obtained in ISZ20 and propagate the uncertainty in the integral for the growth factor. In Figure 3 we compare our results for $f(z)$ with ISZ20 and another previous analysis of BOSS and eBOSS galaxies [50] which assume $\Lambda$CDM. As can be seen, our model-independent results favor smaller values for $f$, which is in qualitative agreement with a recent analysis of the Planck 2018 data in a flat $\gamma\Lambda$CDM model, in which the growth-rate index $\gamma$, defined as $f\simeq\Omega_{m}(z)^{\gamma}$, was left free to vary [51].
[…]
@@ 5 Comparison of distance measurements
so the comparison of the $H_{0}r_{d}$-dependent quantities might be misleading. $F_{AP}$, on the other hand, is independent of $H_{0}r_{d}$. We also include the measurements obtained in ISZ20 with the $\alpha$-analysis, which is a more model-independent analysis of the BOSS data. In this section we always use the mean values and symmetric 68% percentiles.
@@ 6 Conclusions and outlook
There are of course many ways to extend our results in the future. In terms of data, the hexadecapole can be included in the analysis next to monopole and quadrupole. One can also include the bispectrum [56], which is of the same order as the one-loop power spectrum in the non-linear perturbation theory [17]. The BOSS SGC data can also be included, which would increase the effective volume of the sample by around 37% [14]. In terms of methodology, the growth rate $f(z)$ could be generalized to be scale-dependent, and/or one could try using a different number of $k$-bins. One could also use uninformative priors for the dimensionless distance $L_{A}$ to let it be constrained solely by the LSS data. Another possibility is to perform a combined analysis of both redshift bins, since in FreePower they share the same linear $P(k)$ parameters. In terms of theory, our analysis can be expanded by including deviations from the EdS kernels [57, 58, 59]. It can also be extended to include the peculiar-velocity power spectrum [27, 60, 61]. We note that a similar model-independent analysis can be carried out to investigate possible primordial non-Gaussianity (PNG) [16]. We plan to pursue these lines of investigation in future work.
@@ Acknowledgements
We thank Sandro Vitenti for help with the MCMC convergence analysis and Massimo Pietroni and Marco Marinucci for several discussions. MQ is supported by the Brazilian research agencies Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro (FAPERJ) project E-26/201.237/2022, CNPq (Conselho Nacional de Desenvolvimento Científico e Tecnológico) and CAPES. LA acknowledges support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC 2181/1 - 390900948 (the Heidelberg STRUCTURES Excellence Cluster) and under project 456622116. We acknowledge the use of the computational resources of the joint CHE / Milliways cluster, supported by a FAPERJ grant E-26/210.130/2023. This study was financed in part by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brasil (CAPES) - Finance Code 001. We acknowledge support from the CAPES-DAAD bilateral project “Data Analysis and Model Testing in the Era of Precision Cosmology”.
@@ Appendix A Correlations and convergence
The correlation matrices among all 17 parameters are depicted in Figure 6, for both redshift bins. As discussed in the text, most terms are small. The bias parameters have stronger correlations among themselves for the lower redshift bin. The different $P(k)$ bins have strong positive correlations with their immediate neighbors, except when tight priors were used (this was the case for $P_{1}$ for both $z_{\rm eff}$ values, and for $P_{7}$ and $P_{8}$ for $z_{\rm eff}=0.38$). The data-driven cosmological parameters $f$ and $E$ are positively correlated among themselves and anti-correlated with $b_{1}$ and $b_{\mathcal{G}_{2}}$.
We also display the integrated autocorrelation times ($\tau$) in Figure 7. We compute $\tau$ for each walker individually and these estimates are averaged afterwards [37]. The MCMC shows clear signs of convergence as the estimates for the $\tau$ start crossing the dashed line.

### v2

@@ 2 The FreePower model-independent approach
Previous tests indicate that neither the exact spacing nor the number of the $k$-bins has a strong influence on the results on the other parameters [38]. The linear matter power spectrum is constructed with an interpolation between these varied parameters. Because of that, the linear matter power spectrum is quite smooth. Therefore, the coupling between the large-scale IR modes and the mildly non-linear modes is rather weak [10] and the effect of IR resummation is in general small. In this approach we do not disentangle the growth factor $G(z)$ from the linear power spectrum, so we in fact measure $P_{L}(k,z)=G(z)^{2}P_{L}(k)$ at each redshift bin.
@@ 2.2 Data sets from BOSS
linear growth rate $f$ ${\cal U}[0,1]$ ${\cal U}[0,1.5]$
@@
dim.less angular diameter distance $L_{A}$ $\mathcal{N}(0.25,0.008)$ $\mathcal{N}(0.322,0.010)$
linear bias $b_{1}$ $\mathcal{N}(1.67,0.72)$ $\mathcal{N}(2.28,1.04)$
${\cal G}_{2}$ bias $b_{\mathcal{G}_{2}}$ $\mathcal{N}(0.11,1.36)$ $\mathcal{N}(0.6,2.8)$
quadratic bias $b_{2}$ $\mathcal{N}(-2.3,2.8)$ $\mathcal{N}(-2.7,5.2)$
@@
counterterm (monopole) $c_{0}/[{\rm Mpc}/h]^{2}$ $\mathcal{N}(-18.5,95.7)$ $\mathcal{N}(31.5,116)$
counterterm (quadrupole) $c_{2}/[{\rm Mpc}/h]^{2}$ $\mathcal{N}(14.7,72)$ $\mathcal{N}(-43.3,192)$
@@
$P_{7}/$$[{\rm Mpc}/h]^{3}$ $\mathcal{N}(375,500)$ $\mathcal{N}(375,500)$
$P_{8}/$$[{\rm Mpc}/h]^{3}$ $\mathcal{N}(100,200)$ $\mathcal{N}(100,200)$
@@
For the bias parameters, we employed Gaussian priors using as reference the final NGC posterior results from ISZ20. We took their errorbars, propagated the uncertainty in the power spectrum amplitude $A$ since they multiply all their bias parameters by $(A/A_{\rm Planck})^{1/2}$, symmetrized the results and multiplied them by a factor of four. We used the final results for the mean and standard deviation to construct our low-$z$ uncorrelated Gaussian priors. We also did a similar procedure for the counterterm parameters, but only multiplied the results by a factor of two. We thus allow a four times larger uncertainty of ISZ20 in each bias and twice as large for the counterterm parameters as our prior. This resulted in posteriors which were never driven by the priors themselves. I.e., these priors are weakly informative. The Gaussian prior on $L_{A}(z)$ instead is informative and is centered around its Planck 2018 $\Lambda$CDM value (i.e. assuming $\Omega_{m0}=0.31$ so that $L_{A}=0.25$ for the low-$z$ and 0.322 for the high-$z$ sample) with a $3\%$ standard deviation. For $z\leq 1$ current constraints from Pantheon+ [46] are already at this level.
@@ 3 MCMC results
We ran emcee using 256 walkers initialized in a small volume in parameter space. We ran, for each walker, 800k steps, for a total of 200 million chain points for each low-$z$ and high-$z$ samples. Due to the complexity of the 1-loop calculations and the amount of MCMC steps needed, the total computational time used for both redshift bin runs was large, around 100k CPU-hours. This heavy computation cost was the main reason for our choice of a small number of $k$-bins, limiting somewhat the dimensionality of our parameter space. We analyzed the results using both the Python package getdist [47] and our own codes. We compared our final results with those obtained with smaller chains with 400k steps to confirm the results we found were robust, which as we will discuss is in tension with $\Lambda$CDM CMB results. Indeed, the results with half the chain steps change little, which gives us further confidence in the convergence.
@@
From the $P(k)$ linear bins we can also obtain the amplitude $\sigma_{8}$ by interpolating and integrating over a spherical window function. We find $\sigma_{8}$ to be $0.89\pm 0.08$ ($0.82\pm 0.08$) for the $z_{\rm eff}=0.38$ ($0.61$) redshift bin. Note, however, that due to the small number of $k$-bins here considered, the value of $\sigma_{8}$ will inevitably depend on the chosen interpolation scheme. For a standard $\Lambda$CDM-like $P(k)$ binned in the same way as our spectra, our tests show that our chosen interpolation can recover the fiducial $\sigma_{8}$ to within 0.02. Our results are around $1\sigma$ higher than Planck 2018. They also change by less than 2% when discarding the first two prior-dominated $k$ bins.
Table 2 summarizes the marginalized constraints in each cosmological and bias parameter, whereas Table 3 show the results for all 8 $k$ bins of $P(k)$. Both tables also show the best-fit Planck 2018 values assuming flat $\Lambda$CDM.
The correlations $\rho$ among the 17 parameters are in most cases small. The average off-diagonal absolute value of $\rho$ is around 0.16 for both redshift bins. We list here some particular values for $z_{\rm eff}=0.38$ (0.61): $\rho(f,E)=0.59$ ($0.73$), $\rho(E,L_{A})=-0.31$ ($-0.31$), $\rho(E,b_{1})=-0.32$ ($-0.57$), $\rho(c_{0},c_{2})=0.75$ ($0.44$). There is some large correlations between individual $k$-bins of $P(k)$ and some bias parameters. But both $f$ or $E$ show only small or moderate correlations with $P(k)$ and the bias parameters. The full correlation matrices for both bins are shown in Appendix A.
@@
$f$ $0.64^{+0.13}_{-0.11}$ $0.61^{+0.16}_{-0.11}$ 0.716 0.792
$E$ $1.054^{+0.079}_{-0.048}$ $1.27\pm{0.11}$ 1.23 1.41
$L_{A}$ $0.250\pm 0.007$ $0.326\pm 0.010$ 0.250 0.323
$b_{1}$ $2.12^{+0.19}_{-0.18}$ $2.15\pm 0.22$
$b_{2}$ $-1.3^{+1.4}_{-1.9}$ $-2.2^{+5.2}_{-1.3}$
$b_{\mathcal{G}_{2}}$ $0.28^{+0.38}_{-0.29}$ $0.09^{+0.67}_{-0.81}$
$c_{0}/[{\rm Mpc}/h]^{2}$ $9^{+31}_{-51}$ $28^{+38}_{-40}$
$c_{2}/[{\rm Mpc}/h]^{2}$ $46^{+39}_{-46}$ $64^{+59}_{-37}$
$s$ $0.18^{+0.49}_{-0.45}$ $0.07^{+0.42}_{-0.44}$
$\sigma_{8}$ $0.89\pm 0.08$ $0.82\pm 0.08$ $0.811\pm 0.006$
Table 2: Main results for the two redshift bins. We show best-fits with 68.3% CI. The last column shows the expected values for $\Lambda$CDM with Planck 2018 values [48]. Note that $\sigma_{8}$ is a derived parameter from the $P(k)$ bins in Table 3.
@@
Table 3: Best-fit and 68.3% CI for each $k$-bin of the linear power spectrum (at $z=0$), in units of $({\rm Mpc}/h)^{3}$. The $k$ values represent the center of each $k$-bin. The model-independent constraints on $P_{i}(z)$ were converted to $z=0$ values for comparison using the $\Lambda$CDM growth factor $G(z)=0.817\hskip 5.69054pt(0.728)$ for $z=0.38\hskip 5.69054pt(0.61)$. The last row reports the corresponding Planck 2018 $\Lambda$CDM values as reference values.
We tested the convergence of our MCMC using primarly the auto-correlation time ($\tau$) estimates [37]. A large ratio between the number of steps and $\tau$ indicate that the chain ran for long enough to have a large effective number of independent steps. Ideally $\tau$ should converge to an asymptotic value for high enough number of steps. We found that around 400k steps with each of 256 walkers were needed for reaching convergence. More details are provided in Appendix A. We also performed the Gelman-Rubin test [49]. This simple test is not the most reliable, specially using an ensemble sampler like emcee since in this approach the walkers are not independent. Nevertheless it gives another handle on convergence which is independent of the auto-correlation time analysis. For the cosmological parameters $f$, $E$ and $L_{A}$, we get $(\hat{R})^{(1/2)}=\{1.028,1.044,1.016\}$ for the low-$z$ run, and $(\hat{R})^{(1/2)}=\{1.035,1.042,1.018\}$ for the high-$z$ run. It is commonly assumed as a threshold for convergence the value $(\hat{R}_{\rm max})^{(1/2)}=1.1$. This agrees qualitatively with the auto-correlation time results in Appendix A, which indicate reasonable convergence for the chains.
@@ 4 Comparison with previous analyses of BOSS data
A comparison of the inferred cosmological parameters with these works is given in Table 4. In this Table we quote mean values and 68% percentiles using equal-tailed statistics, instead of highest-density values, since those are the ones reported in ISZ20 and PI22. For the low-$z$ survey we find $f=0.66\pm 0.12$ and $E=1.07\pm 0.06$, in reasonable agreement with $f=0.74\pm 0.08$, $E=1.14\pm 0.05$ from ISZ20. FreePower measures directly $f(z)$, while conventional analysis measure instead $f(z)\sigma_{8}(z)$. We therefore convert the latter using the values of $\sigma_{8}$ and $\Omega_{m0}$ obtained in ISZ20 and propagate the uncertainty in the integral for the growth factor.
In Figure 3 we compare our results for $f(z)$ with ISZ20 and another previous analysis of BOSS and eBOSS galaxies [50] which assume $\Lambda$CDM. As can be seen, our model-independent results favor smaller values for $f$, which is in qualitative agreement with a recent analysis of the Planck 2018 data in a flat $\gamma\Lambda$CDM model, in which the growth-rate index $\gamma$, defined as $f\simeq\Omega_{m}(z)^{\gamma}$, was left free to vary [51].
[…]
@@ 5 Comparison of distance measurements
so the comparison of the $H_{0}r_{d}$-dependent quantities might be misleading. $F_{AP}$, on the other hand, is independent of $H_{0}r_{d}$. We also include the measurements obtained in ISZ20 with the $\alpha$-analysis, which is a more model-independent analysis of the BOSS data. In this section we always use the mean values and symmetric 68% percentiles instead of highest-density values (for $L_{A}$ both coincide).
@@ 6 Conclusions and outlook
There are of course many ways to extend our results in the future. In terms of data, the hexadecapole can be included in the analysis next to monopole and quadrupole. One can also include the bispectrum [56], which is of the same order as the one-loop power spectrum in the non-linear perturbation theory [17]. The BOSS SGC data can also be included, which would increase the effective volume of the sample by around 37% [14]. In terms of methodology, the growth rate $f(z)$ could be generalized to be scale-dependent, and/or one could try using a larger number of $k$-bins. One could also use uninformative priors for the dimensionless distance $L_{A}$ to let it be constrained solely by the LSS data. Another possibility is to perform a combined analysis of both redshift bins, since in FreePower they share the same linear $P(k)$ parameters. In terms of theory, our analysis can be expanded by including deviations from the EdS kernels [57, 58, 59]. It can also be extended to include the peculiar-velocity power spectrum [27, 60, 61]. We note that a similar model-independent analysis can be carried out to investigate possible primordial non-Gaussianity (PNG) [16]. We plan to pursue these lines of investigation in future work.
@@ Acknowledgements
We thank Sandro Vitenti for help with the MCMC convergence analysis, Massimo Pietroni and Marco Marinucci for several discussions and Eoin Colgáin for comments on our manuscript. MQ is supported by the Brazilian research agencies Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro (FAPERJ) project E-26/201.237/2022, CNPq (Conselho Nacional de Desenvolvimento Científico e Tecnológico) and CAPES. LA acknowledges support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC 2181/1 - 390900948 (the Heidelberg STRUCTURES Excellence Cluster) and under project 456622116. We acknowledge the use of the computational resources of the joint CHE / Milliways cluster, supported by a FAPERJ grant E-26/210.130/2023. This study was financed in part by the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brasil (CAPES) - Finance Code 001. We acknowledge support from the CAPES-DAAD bilateral project “Data Analysis and Model Testing in the Era of Precision Cosmology”.
Appendix
@@ Appendix A Correlations and convergence
The correlation matrices among all 17 parameters are depicted in Figure 6, for both redshift bins. As discussed in the text, most terms are small. The bias parameters have stronger correlations among themselves for the lower redshift bin. The different $P(k)$ bins have strong positive correlations with their immediate neighbors, with the exception of $P_{1}$, for which tight priors were used. The data-driven cosmological parameters $f$ and $E$ are positively correlated among themselves and anti-correlated with $b_{1}$ and $b_{\mathcal{G}_{2}}$.
We also display the integrated autocorrelation times ($\tau$) in Figure 7. We compute $\tau$ for each walker individually and these estimates are averaged afterwards [37]. The MCMC shows clear signs of convergence as the estimates for the $\tau$ start crossing the $N/20$ line, where $N$ is the number of MCMC steps. Note that since $\tau$ is of the order $10^{4}$ for all parameters, we thin all our chains by a factor of 100 in our analysis, which cuts some of the noise and allows more robust KDE estimation and thus more robust confidence intervals with getdist.